# Identification of Craving Maps among Marijuana Users via the Analysis of Functional Brain Networks with High-Order Attention Graph Neural Networks

### v4

@@ 4 Conclusion
This study proposes a mixed model (HOGANN), which incorporates a dynamic graph LSTM capable of learning temporal patterns in fMRI time series and integrates a high-order attention mechanism. Moreover, the HOGANN is more accurate than other models for predicting marijuana addiction based on the multigraph of brain networks. By utilizing the HOGANN and Yeo atlas of seven intrinsic networks, we identified the most relevant subnetworks and cognitive regions negatively influenced by persistent marijuana use, revealing that chronic marijuana use adversely affects cognitive control, particularly within the dorsal attention and frontoparietal networks, which are essential for attentional, cognitive control, and higher cognitive functions. Therefore, through the HOGANN, we can more accurately predict craving maps in LM users, identify brain maps associated with long-term cravings, and pinpoint active areas important for analysis. In the future, our study not only deepens our comprehension of the effects of long-term marijuana use on the brain but also underscores the revolutionary potential of AI in analyzing functional brain networks.

### v5

@@ 5 Conclusion
In this study, we propose a mixed model (HOGANN) that incorporates a high-order attention mechanism to enhance the neighboring connection features information of fMRI time series within topological graph structures. The HOGANN demonstrates superior accuracy in classifying multigraphs of brain networks and identifying overlapping craving regions. Furthermore, by utilizing the HOGANN in conjunction with the Yeo atlas of seven intrinsic networks, we successfully validated common significant subnetworks across two datasets in LM groups. By identifying specific network alterations associated with marijuana use, our study contributes to the growing body of evidence supporting a multi-faceted, network-based understanding of addiction. Additionally, we effectively elucidated how different demographic factors in the LM group correlate with varying connectivity patterns in brain network craving regions. These findings provide valuable insights that lead to more targeted interventions aimed at normalizing these network dysfunctions in research and treatment of marijuana withdrawal.